# Statistical Multiport-Network Modeling and Efficient Discrete Optimization of RIS

### v1

@@ III-4 Genetic Algorithm (GA)
Figure 2: Achieved SISO channel gain $|S_{21}|^{2}$ for the considered methods as a function of $M$ (horizontal axis) and $\mu_{\mathrm{n}}$ (panels). Each displayed data point is the average over 1500 realizations.
@@ IV Performance Evaluation
IV-A Achieved Channel Gains
Our results displayed in Fig. 2 reveal that in the regime of weak MC (left panel with $\mu_{\mathrm{n}}=0.01$), all algorithms achieve nearly identical channel gains (for sufficiently high $M$ in some cases), even RR-based CD for $M>N_{\mathrm{S}}$. Hence, as one may expect, the affine approximation is very accurate in the weak-MC regime; indeed, the curves for CASC (dashed) and MNT (continuous) are superposed in this case. As the MC level is increased, the approximation underlying CASC naturally becomes less accurate and the differences between CASC and MNT become apparent (middle and right panel). In the regime of moderate MC (middle panel with $\mu_{\mathrm{n}}=0.5$) the performance gap is already notable, and in the regime of strong MC (right panel with $\mu_{\mathrm{n}}=0.99$) it is large. Across all regimes, we observe that MNT-based CD achieves the largest channel gains without any significant dependence on $M$. Thus, no benefits of intelligently initializing CD are apparent. The channel gains achieved with TABP and GA are almost on a par with those of CD for large values of $M$.
IV-B Computational Complexity
@@
[…]
@@
IV-C Memory Complexity
To characterize the memory usage of each optimization algorithm, we estimate the size of the variables stored at each iteration. CD and TABP have constant memory usage, storing only the current configuration (and gradients for TABP), resulting in a memory cost on the order of $\mathcal{O}(1)$, independent of the number of iterations. In contrast, DS, the ridge regression underlying RR-CD, and GA require storing a large set of configurations (for evaluation, surrogate model training, or population evolution), causing their memory usage to grow linearly with the number of configurations, i.e., $\mathcal{O}(M)$.
@@ V Conclusion
To summarize, we introduced a rigorous statistical approach to MNT-based modeling of RIS-parametrized channels; moreover, we leveraged this approach to comprehensively compare optimization strategies for RIS with MC and limited to 1-bit-programmable elements, relevant to most existing RIS prototypes. The choice of optimization method should be guided by the MC strength and the available computational resources. For weak MC, simple methods like DS or RR-CD are sufficient and lightweight. For higher MC, CD and TABP are more effective, with CD offering the best trade-off between performance, speed, and memory. Both methods achieve comparable channel gains and require a similar number of model evaluations; however, TABP requires additional gradient computations, cannot benefit from the Woodbury identity, and inevitably requires a calibrated model. The cost of intelligent initialization could not be amortized, especially not for strong MC. GA can be competitive under parallel execution with loose resource constraints.
Altogether, we conclude that for RIS with non-negligible MC and 1-bit-programmable elements, CD offers the best compromise between channel gain, computational efficiency, and memory usage. These insights also directly apply to beyond-diagonal RIS (based on the physics-compliant diagonal representation [14]); we expect them to generalize to other reconfigurable wave systems like dynamic metasurface antennas and wave-domain physical neural networks.
Looking forward, we envision an extension of our statistical MNT-based model-parameter generation to broadband scenarios. Therein, spectral correlations must be captured correctly, which can be achieved with the so-called Heidelberg random-matrix-theory approach [36, 37]. This approach can be of interest even in the single-frequency case to capture higher-order correlations between the entries of $\tilde{\mathbf{S}}$; however, Rayleigh fading (only concerned with $\mathbf{H}$ in the absence of an RIS) ignores such correlations. Moreover, we plan to extend the presented analysis to account for overhead associated with the estimation of the model parameters.

### v2

@@ IV Performance Evaluation
Figure 2: SISO channel gain $|S_{21}|^{2}$ for the considered methods as a function of $M$ (horizontal axis) and $\mu_{\mathrm{n}}$ (panels), averaged over 1500 realizations.
Achieved Channel Gains: Our results displayed in Fig. 2 reveal that in the regime of weak MC (left panel; $\mu_{\mathrm{n}}=0.01$), all algorithms achieve nearly identical channel gains (for sufficiently high $M$ in some cases), even RR-based CD for $M>N_{\mathrm{S}}$. Hence, as one may expect, the affine approximation is very accurate in the weak-MC regime; indeed, the curves for CASC (dashed) and MNT (continuous) are superposed in this case. As the MC level is increased, the approximation underlying CASC naturally becomes less accurate and the differences between CASC and MNT become apparent (middle and right panel). In the regime of moderate MC (middle panel; $\mu_{\mathrm{n}}=0.5$) the performance gap is already notable, and in the regime of strong MC (right panel; $\mu_{\mathrm{n}}=0.99$) it is large. Across all regimes, we observe that MNT-based CD achieves the largest channel gains without significant dependence on $M$. Thus, no benefits of intelligently initializing CD are apparent. The channel gains of TABP and GA are almost on a par with those of CD for large $M$.
@@
Computational Complexity: We assess the temporal complexity of each optimization algorithm based on the number of evaluations of the model (irrespective of the model’s fidelity) required for convergence or to reach a fixed iteration limit. The number of evaluations equals the dictionary size $M$ for DS, and $10M$ for GA with 10 generations of population size $M$. TABP and CD typically require $\mathcal{O}(N_{\mathrm{S}})$ model evaluations, as seen in Fig. 3. We also observe that better initializations (based on larger $M$) reduce the temporal complexity of TABP and CD by accelerating their convergence, especially in regimes with lower MC strength. For strong MC, the benefits of a good initialization in terms of the convergence speed are small; in fact, the reduction in the number of model evaluations is smaller than $M$ in this case. We also observe that generally the benefits of initialization for the convergence speed saturate when $M>N_{\mathrm{S}}$. Finally, we see in Fig. 3 that TABP and CD generally converge faster for lower MC strengths.
@@
Memory Complexity: To characterize the algorithms’ memory usage, we estimate the size of the variables stored at each iteration. The memory usage of CD and TABP is $\mathcal{O}(N_{\mathrm{S}})$: it only depends on $N_{\mathrm{S}}$ because they only store the current configuration (and gradients for TABP). In contrast, DS, the ridge regression underlying RR-CD, and GA require storing a large set of configurations (for evaluation, surrogate model training, or population evolution); their memory usage grows linearly with the number of configurations, as $\mathcal{O}(MN_{\mathrm{S}})$.
@@ V Conclusion
To summarize, we introduced a simple statistical approach to MNT-based modeling of RIS-parametrized channels and leveraged it to comprehensively compare optimization strategies for RIS with MC and limited to 1-bit-programmable elements, relevant to most existing RIS prototypes. The choice of optimization method should be guided by the MC strength and the available computational resources. For weak MC, simple methods like DS or RR-CD are sufficient and lightweight. For higher MC, CD and TABP are more effective, with CD offering the best trade-off between performance, speed, and memory. Both methods achieve comparable channel gains and require a similar number of model evaluations; however, TABP requires additional gradient computations, cannot benefit from the Woodbury identity, and inevitably requires a calibrated model. The cost of intelligent initialization could not be amortized, especially not for strong MC. GA can be competitive under parallel execution with loose resource constraints.
Altogether, we conclude that for RIS with non-negligible MC and 1-bit-programmable elements, CD offers the best compromise between channel gain, computational efficiency, and memory usage. We expect our insights to generalize to other reconfigurable wave systems (e.g., dynamic metasurface antennas and wave-domain physical neural networks).
Looking forward, we envision an extension of our statistical MNT-based model-parameter generation to broadband scenarios. Therein, spectral correlations must be captured correctly, which can be achieved with the so-called Heidelberg random-matrix-theory approach [35, 36]. This approach can be of interest even in the single-frequency case to capture higher-order correlations between the entries of $\tilde{\mathbf{S}}$; however, Rayleigh fading (only concerned with $\mathbf{H}$ in the absence of an RIS) ignores such correlations. Moreover, the presented analysis can be extended to account for overhead associated with estimating the model parameters.